A 24-volt battery is connected in parallel to a 4-ohm and a 12-ohm resistor.

a. What is ΔV4Ω?

b. How much energy is lost by 1 C of charge as it flows through the 12-ohm resistor?

c. How many coulombs of charge flow through the battery each second? Hints: 1 volt = 1 J/C; 1 ampere = 1 C/s.
Answer #1

42 12 h @ 24V First we need I in circuit RN = 4+12=1662 I = 24 = 22 Amb AV across 48 is V IR = 3 x 4 = 6V as giren IJ Joule oenergy lost E = VX C E = 18x1 =18 J & current drawn from cell lampics 3 Amp = 3/24/2 so, 3 c cell every in flowing through second